reduce the equation


∇²ψ + [k² + f(ρ) + (1/ρ²)g(φ) + h(z)]ψ = 0


to a set of ODEs by the method of separation of variables.


1
Expert's answer
2021-12-01T11:06:22-0500

"\nabla^{2} \psi+\left[k^{2}+f(\rho)+\left(\frac{1}{\rho^{2}}\right) g(\varphi)+h(z)\right] \psi=0"

"\frac{1}{\rho} \frac{\partial}{\partial \rho}\left[\rho \frac{\partial \psi}{\partial \rho}\right]+\frac{1}{\rho^{2} \frac{\partial^{2} \psi}{\partial \varphi^{2}}}+\frac{\partial \psi}{\partial \beta^{2}}+\left[k^{2}+f(\rho)+\left(\frac{1}{\rho^{2}}\right) g(\varphi)+h(z)\right] \psi=0"

"\frac{\partial^{2} \varphi}{\partial^{2}}+\frac{1}{\rho} \frac{\partial \psi}{\partial \rho}+\frac{1}{\rho^{2}} \frac{\partial^{2} \psi}{\partial \varphi^{2}}+\frac{\partial^{2} \varphi}{\partial \underline{ \beta}^{2}}+\left[k^{2}+f(\rho)+\left(\frac{1}{\rho^{2}}\right) g(\varphi)+h(z)\right] \psi=0"

Method of separation of variable

"\psi(\rho, Q, z)=R(\rho) H(\varphi) Z(\beta)"

"R^{\prime \prime} H z+\frac{1}{\rho} R^{\prime} H z+\frac{1}{\rho^{2}} R H^{\prime \prime} z+R H Z^{\prime \prime}+

\left[k^{2}+f(\rho)+\left(\frac{1}{\rho^{3}}\right) g(\varphi)+h(z)\right] R H z=0"

"\frac{R^{\prime \prime}}{R}+\frac{1}{\rho} \frac{R^{\prime}}{R}+\frac{1}{\rho^{2}} \frac{H^{\prime \prime}}{H}+\frac{Z^{\prime \prime}}{Z}+\left[k^{2}+f(\rho)+\left(\frac{1}{\rho^{2}}\right) g(\varphi)+h(z)\right]=0"

"\frac{R^{\prime \prime}}{R}+\frac{1}{\rho} \frac{R^{\prime}}{R}+\frac{1}{\rho^{2}} \frac{H^{\prime \prime}}{H}+\left[k^{2}+f(\rho)+\left(\frac{1}{\rho^{2}}\right) g(\varphi)\right]=\frac{-Z^{\prime \prime}}{z}-h(z)

=1"

Then "\quad-\frac{z^{\prime \prime}}{z}-R(\xi)=\lambda \Rightarrow z^{\prime \prime}+(\lambda+R(3)) z=0" and "\frac{R^{\prime \prime}}{R}+\frac{1}{\rho} \frac{R^{\prime}}{R}+\frac{1}{\rho^{2}} \frac{H^{\prime \prime}}{H}+\left[k^{2}+f(\rho)+\left(\frac{1}{\rho^{\prime}}\right) g(\varphi)\right]=\lambda"

"\begin{aligned}
&\Rightarrow \quad \frac{\rho^{2} R^{\prime \prime}}{R}+\frac{\rho R^{\prime}}{R}+\frac{H^{\prime \prime}}{H}+\left(K^{2} \rho^{2}+\rho^{2} f(\rho)+g(\varphi)\right)=\lambda \rho^{2}\\
&\Rightarrow \frac{\rho^{2} R^{\prime \prime}}{R}+\frac{\rho R^{\prime}}{R}+K^{2} \rho^{2}+\rho^{2} f(\rho)-\lambda \rho^{2}=\frac{-H^{\prime \prime}}{H}-g(\varphi)=l \\
&\text { Then } \quad \frac{-H^{\prime \prime}}{H}-g(\varphi)=l \Rightarrow H^{\prime \prime}+(l+g(\varphi)) H=0\\
&\text { and } \frac{\rho^{2} R^{\prime \prime}}{R}+\frac{\rho R^{\prime}}{R}+R^{2} \rho^{2}+\rho^{2} f(\rho)-\lambda \rho^{2}=l\\
&\Rightarrow \quad \rho^{2} R^{\prime \prime}+\rho R^{\prime}+\left(K^{2} \rho^{2}+\rho^{2} f(\rho)-\lambda \rho^{2}-\ell\right) R=0\\
&\text { set of ODE's }\\
&\begin{gathered}
z^{\prime \prime}+(\lambda+R(z))z=0 \\
H^{\prime \prime}+(l+g(Q)) H=0 \\
\rho^{2} R^{\prime \prime}+\rho R^{1}+\left(K^{2} \rho^{2}+\rho^{2} f(\rho)-\lambda \rho^{2}-l\right) R=0
\end{gathered}\\
\end{aligned}"